Exercise 17-6 Assigning overhead costs using the plantwide rate and departmental rate methods LO P1, P2 Laval produces lampsExercise 17-6 Assigning overhead costs using the plantwide rate and departmental rate methods LO P1, P2 Laval produces lampsExercise 17-6 Assigning overhead costs using the plantwide rate and departmental rate methods LO P1, P2 Laval produces lampsExercise 17-6 Assigning overhead costs using the plantwide rate and departmental rate methods LO P1, P2 Laval produces lamps
Answer #1

Solution 1:

Total direct labor hours for both departments = 210,000 + 340,000 = 550,000 DLH

Total overhead cost = $350,000 + $435,000 = $785,000

Plantwide overhead rate = Estimated overhead costs / estimated direct labor hours  = $785,000 / 550,000 = $1.43 per DLH

Solution 2

Direct materials $265,000
Direct labor -Fabricating $252,000
Direct labor - Assembly $354,200
Overhead (1.43 x 24400) $34892
Total manufacturing cost $906,092
Units produced 25000
Manufacturing cost per unit $36.24

Solution 3:

Computation of departmental overhead rate
Particulars Fabricating    Assembly
Overhead cost of deparment    $350,000 $435,000
Allocation base MH DLH
Total machine hours 169,000
Total direct labor hours 340,000
Overhead rate of department $2.07 MH $1.28 DLH   

Solution 4:

Direct materials $265000
Direct labor
Fabricating $252000
Assembly $354200
606,200
Overhead
Fabricating ($2.07* 15100) 31272
Assembly ($1.28* 15400) 19703
50975
Total manufacturing cost $922,175
Units produced 25000
Manufacturing cost per unit $36.89